Q3. A flat rectangular piece of aluminum has a perimeter of 62 inches. The length is 15 inches longer than the width. Find the width.
a. 31 inches
b. 38 inches
c. 8 inches
d. 23 inches

Answer by fractalier(6550)

Call the length and width, L and W. Then we have
Perimeter P = 2L + 2W = 62 and
L = W + 15
Now substitute this into the first equation and get
2(W + 15) + 2W = 62
2W + 30 + 2W = 62
4W + 30 = 62
4W = 32
W = 8 inches
Choice C.